# RAD: Towards Trustworthy Retrieval-Augmented Multi-modal Clinical Diagnosis

### v1

@@ C.1.3 Details of Harvard-FairVLMed and SkinCAP
The Harvard-FairVLMed dataset [40], sourced from the Department of Ophthalmology at Harvard Medical School, contains 10,000 multimodal samples (7,000 train, 1,000 val, 2,000 test) with paired clinical notes, diagnostic labels, and detailed demographic attributes (race, gender, ethnicity, language). The dataset is publicly available under the CC BY-NC-ND 4.0 license at Github [^3] https://github.com/Harvard-Ophthalmology-AI-Lab/FairCLIP. We directly used the original dataset.
SkinCAP is a multimodal dermatology dataset containing 4,000 expert-annotated skin disease images with rich natural language descriptions [75]. The dataset combines cases from diverse dermatology image datasets, all annotated by board-certified dermatologists to ensure clinical accuracy. It is publicly available under an open license at HuggingFace [^4] https://huggingface.co/datasets/joshuachou/SkinCAP. To address class imbalance, we removed tail categories with too few positive samples, resulting in a filtered dataset of 2,526 samples with 50 disease labels. The final dataset was partitioned into a 4:1 train-test split.
@@ C.2 Benchmarking MIMIC-ICD53
To further evaluate the quality of our constructed dataset MIMIC-ICD53, we employed various unimodal methods to train and test its performance. For the visual modality, we selected ResNet-18, ResNet-50 [21], ViT [15], and Swin Transformer [38], along with two SOTA CXR-specific pretrained models, UniChest [14] and RadDino [44] as the baselines. Based on both computational efficiency and data leakage prevention considerations, we ultimately designated ResNet-50 as the standard visual backbone for the main experiments. ViT is also investigated in ablation studies in Section C.6.
For electronic health record (EHR) data, we first leveraged its inherent tabular structure by treating each EHR attribute as an input dimension, with corresponding numerical values assigned to their respective dimensions. We experimented with MLP and a SOTA tabular data process method TabFPN [23], but both exhibited suboptimal performance. Consequently, we reformatted the EHR data into natural language text using the following template: "Laboratory values within the 4–7 range indicate normal levels, values 0–3 suggest clinically low levels, and values 8–10 denote elevated levels. The current panel includes [ATTRIBUTE] with the discretized value of [VALUE]…". We then evaluated the reformatted EHR data using classic backbone ClinicalBERT [54], BioClinicalBERT [2], and PubMedBERT [18]. We also included natural and medical LLMs LLaMa [53] and MMed-S [57]. Specifically, we replace the last layer of LLMs with a classification head to adapt to the text classification task. Due to the high consumption of computing resources, we only use LoRA [24] to fine-tune the LLM-based models. We further conducted experiments using the same text encoders for the report modality alone and reports combined with EHR data. As shown in Table 6, unimodal performance analysis reveals that the report modality achieves the highest diagnostic results, followed by the EHR modality. Combining the two modalities in text form (Report+EHR) can bring significant performance gains. For the visual modality, the performance gap between different backbone architectures is relatively small. While for the text-based modality, LLM-based models generally outperform BERT-based models. Among the BERT-based models, ClinicalBERT consistently achieves the best performance. Considering model size and practicality, we selected ClinicalBERT as the default text encoder in our RAD framework.
@@ C.6 Ablation Study
Ablation on different backbones.
Ablation on LLM refinement of the retrieved knowledge.
@@
C.7 Hyper-parameter analysis

### v2

@@ C.1.3 Details of Harvard-FairVLMed and SkinCAP
The Harvard-FairVLMed dataset [luo2024fairclip], sourced from the Department of Ophthalmology at Harvard Medical School, contains 10,000 multimodal samples (7,000 train, 1,000 val, 2,000 test) with paired clinical notes, diagnostic labels, and detailed demographic attributes (race, gender, ethnicity, language). The dataset is publicly available under the CC BY-NC-ND 4.0 license at Github [^3] https://github.com/Harvard-Ophthalmology-AI-Lab/FairCLIP. We directly used the original dataset.
SkinCAP is a multimodal dermatology dataset containing 4,000 expert-annotated skin disease images with rich natural language descriptions [zhou2024skincap]. The dataset combines cases from diverse dermatology image datasets, all annotated by board-certified dermatologists to ensure clinical accuracy. It is publicly available under an open license at HuggingFace [^4] https://huggingface.co/datasets/joshuachou/SkinCAP. To address class imbalance, we removed tail categories with too few positive samples, resulting in a filtered dataset of 2,526 samples with 50 disease labels. The final dataset was partitioned into a 4:1 train-test split.
@@ C.2 Benchmarking MIMIC-ICD53
To further evaluate the quality of our constructed dataset MIMIC-ICD53, we employed various unimodal methods to train and test its performance. For the visual modality, we selected ResNet-18, ResNet-50 [he2016deep], ViT [dosovitskiy2020image], and Swin Transformer [liu2021swin], along with two SOTA CXR-specific pretrained models, UniChest [dai2024unichest] and RadDino [perez2025exploring] as the baselines. Based on both computational efficiency and data leakage prevention considerations, we ultimately designated ResNet-50 as the standard visual backbone for the main experiments. ViT is also investigated in ablation studies in Section C.6.
For electronic health record (EHR) data, we first leveraged its inherent tabular structure by treating each EHR attribute as an input dimension, with corresponding numerical values assigned to their respective dimensions. We experimented with MLP and a SOTA tabular data process method TabFPN [hollmann2025accurate], but both exhibited suboptimal performance. Consequently, we reformatted the EHR data into natural language text using the following template: "Laboratory values within the 4–7 range indicate normal levels, values 0–3 suggest clinically low levels, and values 8–10 denote elevated levels. The current panel includes [ATTRIBUTE] with the discretized value of [VALUE]…". We then evaluated the reformatted EHR data using classic backbone ClinicalBERT [wang2023optimized], BioClinicalBERT [alsentzer2019publicly], and PubMedBERT [gu2021domain]. We also included natural and medical LLMs LLaMa [touvron2023llama] and MMed-S [wu2025towards]. Specifically, we replace the last layer of LLMs with a classification head to adapt to the text classification task. Due to the high consumption of computing resources, we only use LoRA [hu2022lora] to fine-tune the LLM-based models. We further conducted experiments using the same text encoders for the report modality alone and reports combined with EHR data. As shown in Table 6, unimodal performance analysis reveals that the report modality achieves the highest diagnostic results, followed by the EHR modality. Combining the two modalities in text form (Report+EHR) can bring significant performance gains. For the visual modality, the performance gap between different backbone architectures is relatively small. While for the text-based modality, LLM-based models generally outperform BERT-based models. Among the BERT-based models, ClinicalBERT consistently achieves the best performance. Considering model size and practicality, we selected ClinicalBERT as the default text encoder in our RAD framework.
@@ C.6 Ablation Study
Ablation on different backbones.
Ablation on LLM refinement of the retrieved knowledge.
@@
C.7 Cost Analysis.
C.8 Hyper-parameter analysis